a birthday party balloon is filled with 0.045 grams of helium gas. if the balloon takes up 750.ml of space and has a pressure of 325 mmhg. what is the teperature inside the balloon
Answer #1

mass fo helium = 0.045 grams

molar mass of helium= 4 grams/mole

volume = 750 ml= 0.750 L

pressure = 325 mmhg = 325/760 = 0.4276 atm

Gas constant = R = 0.0821 L-atm/mole-k

Ideal gas equation              PV=nRT                  But n= mass/molar mass = 0.045/4 = 0.01125 moles

                                  T = PV/nR

                                   T = 0.4276 x 0.750/ 0.01125 x0.0821

                                   T = 0.3207/0.000924

                                T = 340.08 K

Tempature present inside of the balllon = T= 340.08 K.